OpenStudy (anonymous):

2x10 to the third power x 4x10 to the fifth power is this this the answer:2,000 x400,000...it says to multiply and write the answer without exponets

OpenStudy (mathstudent55):

Use ^ for exponent. For example, 10 to the thrid power is written 10^3. 10^1 = 10 10^2 = 10 x 10 = 100 10^3 = 10 x 10 x 10 = 1,000 So 2 x 10^3 = 2 x 1,000 = 2000 and 4 x 10^5 = 4 x 100,000 = 400,000 How do you multiply numbers with exponents?

OpenStudy (anonymous):

Remember Order of Operations....Parenthesis..Exponents...Multiplication..Division...Addition...Subtraction...working from left to right

OpenStudy (anonymous):

what would be the area of a triangle if the base is 1& 3/4 and the height is 4/5. The area of a triangle equals half the product of its base and height. What would the area be? Can you show me how to figure out?

OpenStudy (anonymous):

solve the system of linear equations by graphing. 3x-y=0 2x-y=3. type an ordered pair, or is there infinitely many soultions or no solution.